Operations Report – April 14th

Crew 210 Operations Report 14-APR-2019

SOL: 2

Name of person filing report: Shannon Rupert

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run): 11.5 hours last night

Solar— SOC unknown %. Will not be reported during COMMS as I plan to
turn it on as late as possible. I find recording these data on this
report awkward. Is there a better way to report when the generator is
turned off and on and the hours it is on each day?

Diesel Reading – 46 %

Propane Reading – 46 %

Ethanol Free Gasoline – 10 gallons.

Water (auxillary tank) – 0 gallons

Water (static tank) – 200 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – unknown gallons.

Water (loft) – Static to Loft Pump used – yes

Water Meter: not read

Toilet tank emptied: no

Deimos rover used: Still non-operational


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, no

Hours: 50.7

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: No

Hours: 86.5

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: yes

Hours: 56.4

Beginning charge: 100

Ending charge: 97

Currently charging: yes

Curiosity rover used: No

Hours: 88.8

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) None

Reason for use: None were used.

Oil Added? No

ATV Fuel Used: 0 Gals.

# Hours the ATVs were Used today: 0

Notes on ATVs: We are planning a morning trip to Loa tomorrow for two
more batteries and a spark plug for the Honda. Also some battery
screws and bolts. Two of the 350’s are operational: the one that
originally started and the one with the new battery, The Honda just
won’t turn over, even with canned starter. The battery in the 300 was
damaged somehow and needs to be replaced. The remaining 350’s battery
will not hold a charge and also needs to be replaced.

HabCar used and why, where? Yes, to Salt Creek for research. We had
a problem with it not starting once we returned to MDRS and went to
the Quarry. It seems the problem is in the shift column, not the
starter or the battery. We finally got it started and returned to the
Hab only minutes after the two crewmembers who walked back to the Hab
had arrived.

CrewCar used and why, where? No

General notes and comments: Tunnel repairs were almost completed this
morning. The remaining work will be done tomorrow morning.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Still not completely cleaned up.
David will be sending a GreenHab report tonight and from now on.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing tonight,
thank you.

Science Report – April 14th

Sol 2 – Science Summary

Crew 210 sampled vascular plants and lichens from three sites across the MDRS operational area today; the Valley of the Stars, Salt Creek, and the vicinity of the Burpee Dinosaur Quarry. We collected between 15-20 samples, many of which were species not recorded during previous floristic work at the station in 2014. We will press the samples later this evening to preserve them for later identification. Tomorrow, if time permits, we will continue our sampling in the area just around the hab.

Sol Summary – April 14th

Sol 2 – Anywhere is walking distance if you’ve got the time

Before departing MDRS for a day of botanical sampling, crew 210 spent the morning working on social media outreach and (importantly), fueling up on coffee. While team members emptied the hab car water tank (filled last night in Hanksville), other crew members continued to repair the hab tunnels. The full team (five humans and two space dogs) left the station at 11:30.

After refueling the hab car in Hanksville, we drove to the Factory Butte access road and sampled plants and lichens at the Valley of the Stars and at Salt Creek. We returned to Cow Dung Road and went to continue our plant sampling outside of the Burpee Dinosaur Quarry

When we went to return to MDRS the hab car did not initially start, so Mike Irvine and myself hiked back to the hab from the quarry to retrieve the crew car. However, just as we returned to the hab we heard from the rest of the team over the radio that they were able to restart the car and they made their way back to MDRS.

Tomorrow the team will travel to Loa in the morning to retrieve parts for the ATVs, and plan on sampling around the hab (time-permitting) in the afternoon.

Greenhab Report – April 14th

GreenHab Officer: David Murray

Environmental control: Ambient with door open from 10:40 to 19:19

Shade Cloth: 80%

Average Temperature:

Low Temperature: 14 C

High Temperature: 43 C

Relative humidity: 25%

Hours of Supplemental Light: 0

Daily Water Usage for Crops: 13 L

Water in Blue Tank: N/A

Time(s) of Watering: 10:40 and 19:19

Changes to Crops: None.

Narrative: Our crew spent the day at Salt Creek, a few miles away from the station. Before we left, we watered the crops and left the GreenHab door open but the interior temperature still reached 43 C while we were gone. Since outdoor temperatures are likely to increase until the fall, we’ll probably have to run the fan in addition to propping the door open from here on out. I’ll program the fan to operate during the hottest parts of the day and report the schedule tomorrow. There are discrepancies among the thermometers in the GreenHab and a couple of them have problems with their digital displays. They should be replaced.

Harvest: None

Copyright © The Mars Society. All rights reserved. | Main Site